OpenStudy (anonymous):

Abraham currently has a balance of $3,075.21 in an account he has held for 13 years. He opened the account with an initial deposit of $2,152. What is the simple interest rate on the account?

OpenStudy (texaschic101):

I = PRT for R R = I/PT P = 2152, T = 13, I = 3075.21 - 2152 = 923.21 R = I/PT R = 923.21/(2152)(13) R = 923.21/27976 R = .033 put it in percent form = 3.3% check... I = PRT 923.21 = (2152)(.033)(13) 923.21 = 923.208 rounds to 923.21 (correct) The interest rare is 3.3%

OpenStudy (anonymous):

Thanks, a medal for ya

OpenStudy (texaschic101):

no problem.....got a little confused with the interest...finally figured out that you have to do some subtracting :)